100g of ice at 0°C is mixed with 100g of water at 100°C what will be the final temperature of mixture

Grade:

1 Answers

Eshan
5 years ago
Dear student,

Amount of heat required in melting 100g of ice=mL_f=100\times 80cal=8kcal

Let the final temperature of water beT.

From conservation of energy,

Heat required to melt the ice+Heat required to raise the temperature of melted ice to ‘T’=Heat given out by cooling to water at100^{\circ}Cto ‘T’

\implies 8kcal+0.1(1)(T-0)=0.1(1)(100-T)
\implies T=10^{\circ}C